The mean age of six persons is 49 years. The ages of five of these six persons are 59, 38, 38, 51, and 30 years. Find the age of the sixth person.
Answer by Cromlix(4381)

If the mean age of 6 people is 49
then the total age = 6 * 49 = 294
Adding the ages of the 5 = 216
294 - 216 = 78
This is the age of the 6th person.
Hope this helps you.
:-)